ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 111, 1999

Csörnyei, M. Aronszajn null and Gaussian null sets coincide  191

Ellers, E. W., N. Gordeev and M. Herzog Covering numbers for Chevalley groups

339

Giordano, T., I. F. Putnam and C. F. Skau Full groups of Cantor minimal systems

285

Gordeev, N. See E. W. Ellers

Granero, A. S., M. Jiménez Sevilla and J. P. Moreno Geometry of Banach spaces with property β

263

Herzog, M. See E. W. Ellers

Hjorth, G. Orbit cardinals: On the effective cardinalities arising as quotient spaces of the form X/G where G acts on a Polish space X

221

Iwanik, A. and J. Serafin Code length between Markov processes  29

Jiménez Sevilla, M. See A. S. Granero

Kaplan, G. Irreducible disconnected systems in groups  203

Karpenko, N. A. Three theorems on common splitting fields of central simple algebras

125

Kohlberg, E. and A. Neyman A strong law of large numbers for nonexpansive vector-valued stochastic processes

93

Marcus, M. and A. J. Zaslavski On a class of second order variational problems with constraints

1

Moreno, J. P. See A. S. Granero

Müller, E. F. and H.-J. Schneider Quantum homogeneous spaces with faithfully flat module structures

157

Neyman, A. See E. Kohlberg

Putnam, I. F. See T. Giordano

Ran, Z. Bend, break and count  109

Rowen, L. H. and Y. Segev The finite quotients of the multiplicative group of a division algebra of degree 3 are solvable

373

Rudelson, M. Almost orthogonal submatrices of an orthogonal matrix

143

Scherbak, I. and A. Szpirglas Lagrange transformations and duality for corner and flag singularities 77 Schneider, H.-J. See E. F. Müller Seelinger, G. F. Brauer–Severi schemes of finitely generated

algebras

321

Segev, Y. See L. H. Rowen Serafin, J. See A. Iwanik Skau, C. F. See T. Giordano Szpirglas, A. See I. Scherbak

Talagrand, M. Concentration and influences  275

Zaslavski, A. J. See M. Marcus Živaljević, R. T. The Tverberg–Vrećica problem and the

combinatorial geometry on vector bundles 53